OpenStudy (anonymous):

Evaluate the limit \lim_{ x \rightarrow 0 } \frac {\sin(4x) } {9x }

OpenStudy (mimi_x3):

\[\lim_{ x \rightarrow 0 } \frac {\sin(4x) } {9x }\] using l'hopital's rule. \[ \lim_{ x \rightarrow 0 } \frac {4cos(4x) } {9 }= \frac{4}{9}\]